1. An astronaut leaves earth and travels at 0.92c. He spends 15 years exploring space. When he returns, how much time will have passed on Earth?
1
Expert's answer
2021-10-11T16:45:05-0400

The time on Earth is

"t_0=\frac{t}{\sqrt{1-v^2/c^2}}=38.3\text{ years}."